### v1

# DiffAU: Diffusion-Based Ambisonics Upscaling

Abstract
Spatial audio enhances immersion by reproducing 3D sound fields, with Ambisonics offering a scalable format for this purpose. While first-order Ambisonics (FOA) notably facilitates hardware-efficient acquisition and storage of sound fields as compared to high-order Ambisonics (HOA), its low spatial resolution limits realism, highlighting the need for Ambisonics upscaling (AU) as an approach for increasing the order of Ambisonics signals. In this work we propose DiffAU, a cascaded AU method that leverages recent developments in diffusion models combined with novel adaptation to spatial audio to generate 3rd order Ambisonics from FOA. By learning data distributions, DiffAU provides a principled approach that rapidly and reliably reproduces HOA in various settings. Experiments in anechoic conditions with multiple speakers, show strong objective and perceptual performance.
Index Terms—  Spatial Audio, Ambisonics, Diffusion.
1 Introduction
[…]
Several methods have been proposed in the literature for AU. Early approaches are predominantly model-based, and rely on physical assumptions on the sound field. A representative approach applies compressed sensing (CS) techniques for plane wave decomposition (PWD) under the assumption that the sound field is sparse in the source domain [5, 6, 7]. While CS techniques enable upscaling under ideal conditions, such as free-field and low-noise settings, their performance deteriorates when the sparsity assumption does not hold, limiting their applicability to real-world acoustic environments.
To address the limitations of model-based AU, several data-driven methods have been introduced in recent years. Gao et al. [8] proposed a multi-scale convolutional network operating in the frequency domain, incorporating sparse encoding to enhance generalization. While this method demonstrated improved performance over classical counterparts when the sparsity assumption holds, it similarly struggles in scenarios where this assumption fails. Routray et al. [9] designed a multi-stage deep neural network (DNN) where each stage incrementally upsamples by one order using fully connected networks. Despite its novel hierarchical structure, the architecture lacks expressivity and a theoretical basis. More recently, Nawfal et al. [10] employed a waveform-domain encoder-decoder architecture adapted from Conv-TasNet [11], enabling upscaling in a latent space. Although showing significant improvement over FOA, the reported listening test showed some gaps from ideal 3rd order Ambisonics and from HOA. These limitations motivate the search for methods that can further improve the quality of AU. Generative networks, which were not explored for this task to-date, offer a promising solution.
In this work, we propose DiffAU, a novel cascaded framework for AU that integrates the hierarchical structure of Ambisonics orders with the generative capabilities of diffusion models. By treating AU as a structured generative task, DiffAU enables principled upscaling from FOA to HOA (3rd order in this work), through a sequence of intermediate stages, each implemented as a conditional diffusion process. This approach offers key conceptual advantages: it avoids explicit prior assumptions such as source sparsity and provides a probabilistic mechanism for resolving the underdetermined nature of AU. Moreover, the modularity of DiffAU allows flexible extension to arbitrary upscaling ranges and facilitates order-by-order interpretability and training.
Our design begins by formulating the AU problem as conditional sampling using stochastic differential equations, where the goal is to sample HOA coefficients conditioned on lower-order observations. Based on this formulation, we develop a tailored diffusion model for spatial audio, incorporating signal representations aligned with Ambisonics encoding and appropriate transformations. Each diffusion stage is trained independently using denoising score matching, and the full system is realized via a cascaded architecture. Extensive experiments demonstrate that DiffAU systematically outperforms available AU baselines, highlighting the potential of generative diffusion-based techniques for spatial audio applications.
The rest of this paper is organized as follows: Section 2 introduces the signal model and some preliminaries. Section 3 describes DiffAU, while Sections 4-5 present its empirical study and listening test. Section 6 provides concluding remarks.
2 System Model and Preliminaries
2.1 Signal Model and Ambisonics
Consider a sound field composed of $Q$ plane waves with directions of arrival (DOAs) $(\theta_{q},\phi_{q})$, $q\in\{1,\ldots,Q\}$. The $Q\times 1$ vector ${\boldsymbol{s}}(k)$ is the source signals, with each element corresponding to the amplitude of a plane wave at the origin, with $k$ the wave number. The Ambisonics signal of order $N$ due to ${\boldsymbol{s}}(k)$ and the $Q$ plane waves can be written as [12]
${\boldsymbol{a}}_{N}(k)={\boldsymbol{Y}}_{Q}^{H}{\boldsymbol{s}}(k),$ (1)
where ${\boldsymbol{a}}_{N}(k)$ is size $(N+1)^{2}$ and ${\boldsymbol{Y}}_{Q}$ is a $Q\times(N_{a}+1)^{2}$ matrix with elements the spherical harmonics functions [12] at directions $(\theta_{q},\phi_{q}$). This Ambisonics signal can be computed from microphone signals using a spherical array of radius $r$ [12], and to avoid spatial aliasing for $k\cdot r\ll N$, it must hold that $(N+1)^{2}\leq M$, where is the number of microphones in the array.
2.2 Problem Formulation
AU refers to the mapping of a low-order Ambisonics signal of order $N$ into higher-order coefficients to obtain a signal of order $N^{\prime}>N$. This as is formulated an inverse problem
${\boldsymbol{a}}_{N}(k)={\boldsymbol{F}}{\boldsymbol{a}}_{N^{\prime}}(k),$ (2)
where ${\boldsymbol{a}}_{N}(k)$ is an Ambisonics signal of order $N$ with $(N+1)^{2}$ channels, ${\boldsymbol{a}}_{N^{\prime}}(k)$ is an Ambisonics signal of order $N^{\prime}$ with $(N^{\prime}+1)^{2}$ channels, and ${\boldsymbol{F}}\in\mathbb{R}^{(N+1)^{2}\times(N^{\prime}+1)^{2}}$ is a sampling matrix which takes the first $(N+1)^{2}$ channels of ${\boldsymbol{a}}_{N^{\prime}}(k)$.
Since $\boldsymbol{F}$} is a wide matrix, the problem is underdetermined. A straightforward method for tackling it is the least-norm approach; however, as demonstrated in [13], this tends to distribute the energy uniformly across the plane-wave sources, resulting in distortions. Therefore, incorporating prior knowledge is essential to obtain a physically plausible solution. To explain the concept guiding our methodology for learning $p({\boldsymbol{a}}_{N^{\prime}}(k))|{\boldsymbol{a}}_{N}(k))$ introduced in Section 3, we utilize emerging tools based on score-based generative models, reviewed next.
2.3 Preliminaries of SGMs
SGMs [14] are diffusion-based generative models that learn to reverse a noise corruption process. The formulation of [15] casts this process into a continuous-time SDE, providing a unifying framework for SGMs. The forward process is expressed as
$d{\boldsymbol{x}}_{t}=f({\boldsymbol{x}}_{t},t)dt+g(t)d{\boldsymbol{w}},$ (3)
where $f$ is the drift term, $g$ the diffusion coefficient, and $\boldsymbol{w}$} a standard Wiener process. The corresponding reverse-time dynamics [16] are
$d{\boldsymbol{x}}_{t}=[f({\boldsymbol{x}}_{t},t)dt-g(t)^{2}\nabla_{\mathbf{x}}\log p_{\text{data}}(\mathbf{x})]+g(t)d{\boldsymbol{\bar{w}}},$ (4)
with $\boldsymbol{\bar{w}}$} a time-reversed Wiener process. The score function $\nabla_{{\boldsymbol{x}}}\log p_{\text{data}}({\boldsymbol{x}}_{t})$ is approximated by a neural network ${\boldsymbol{s}}({\boldsymbol{x}}_{t},t;{\boldsymbol{\theta}})$. Sampling can then be carried out using Predictor-Corrector samplers [15, Appendix G].
3 Proposed Method
SGMs have been proposed for image super-resolution, .e.g., [17, 18]. However, applying this methodology to AU is not straightforward, and existing image super-resolution methods do not directly transfer to audio. Still, the success of this approach in other domains motivates exploring its adaptation to spatial audio. To that end, we introduce the proposed DiffAU framework.
3.1 DiffAU
To explain how we adapt diffusion-based super resolution to spatial audio, we begin by formulating the diffusion model and SDE. Next, we describe the data representation employed in our diffusion models, and present our proposed DiffAU for AU.
Fig. 1: Schematic illustration of the overall architecture of DiffAU
SDE Formulation for Spatial Audio: The formulation presented in Subsection 2.3 is geared towards sampling from a prior distribution of some variable $\boldsymbol{x}$}, which in our case represents higher-order Ambisonics. However, we are interested in sampling from the posterior distribution $p({\boldsymbol{x}}|{\boldsymbol{y}})$ conditioned on an observation $\boldsymbol{y}$}, e.g., a lower-order Ambisonics. This setting requires a conditional diffusion model which must estimate the gradient $\nabla_{{\boldsymbol{x}}}\log p_{\text{data}}({\boldsymbol{x}}|{\boldsymbol{y}})$. Conditional diffusion can be achieved by concatenating the observation $\boldsymbol{y}$} to the input of the score model [19]. In our proposed method, we adopt this strategy by using a lower order Ambisonics, FOA in our case, as the observation.
The SDE we use is the Variance Exploding (VE) SDE introduced in[15]. In this formulation, the drift and diffusion terms in the forward (3) and reverse (4) SDEs are defined as follows
$f({\boldsymbol{x}}_{t},t)=0,\quad g(t)=\sigma_{\text{min}}\left(\frac{\sigma_{\text{max}}}{\sigma_{\text{min}}}\right)^{t}\sqrt{2\log\left(\frac{\sigma_{\text{max}}}{\sigma_{\text{min}}}\right)}.$ (5)
Here, $\sigma_{\min}$ and $\sigma_{\max}$ denote the minimum and maximum noise levels specified by the scheduler.
Data Representation: The input signals are order-$N$ real Ambisonics with N3D normalization [2] in the time domain ($N=1$ for FOA), represented as a sequence of $(N+1)^{2}\times 1$ vectors $\{{\boldsymbol{a}}_{N}(\tau)\}_{\tau=1}^{{\tau_{\max}}}$, which can be arranged column-wise into a matrix ${\boldsymbol{A}}^{\tau}_{N}$ of size $(N+1)^{2}\times{\tau_{\max}}$. For SGM-based processing, we apply the short-time Fourier transform (STFT) to obtain the time-frequency (TF) representation ${\boldsymbol{A}}^{f}_{N}$, followed by a nonlinear amplitude transformation $\mathcal{H}(x)$ [20, 21] to normalize the heavy-tailed speech amplitudes and ensure consistent DNN input scales:
$\mathcal{H}(x)=\frac{|x|^{\alpha}}{\beta}e^{i\arg(x)},\quad\mathcal{H}^{-1}(x)=\beta|x|^{1/\alpha}e^{i\arg(x)}.$ (6)
Real and imaginary parts are concatenated along the channel dimension to form real-valued inputs. For the output HOA, predicted channels are first combined into a complex signal, transformed with $\mathcal{H}^{-1}$, and then inverted via inverse STFT (ISTFT). The reconstructed channels are finally concatenated with the input lower-order Ambisonics.
Overall Algorithm: DiffAU, whose overall procedure is illustrated in Fig. 1, consists of two cascaded diffusion blocks. Each block is responsible for upscaling the signal by one Ambisonics order, ultimately producing a third-order Ambisonics representation from the FOA. Each block that maps order $N\mapsto N+1$ is conditioned on the current order signal ($(N+1)^{2}$ channels), and predicts the additional $(2N+3)$ channels required to reach the next order. The backbone of our score model is noise-conditioned score-matching network (NSCN++) architecture [15], which is a Unet with progressive growth, whose parameters at the $N$th block are denoted by ${\boldsymbol{\theta}}_{N}$. We denote the score model for each block as $s_{N}\left({\boldsymbol{x}}_{t},{\boldsymbol{y}};{\boldsymbol{\theta}}_{N}\right)$, where $\boldsymbol{y}$} is the Ambisonics signal of order $N$, and ${\boldsymbol{x}}_{t}$ represents the noisy missing channels that need to be predicted to complete the order $N+1$ signal.
For the sampling process, we adopt the predictor corrector framework as described in [15, Appendix G], employing the reverse diffusion method as the predictor [15, Appendix E], which serves as a discretized approximation of the reverse-time SDE in (4). For the corrector, we use the annealed Langevin dynamics approach [15, Alg. 4] with SNR parameter $0.5$. The complete sampling procedure is outlined in Algorithm 1. Since the DNN operates on real-valued signals, we convert the complex inputs into real-valued ones by stacking the real and imaginary parts, while representing the outputs as complex-valued, with the corresponding transformations denoted $\mathcal{R}$ and $\mathcal{C}$, respectively. In our implementation, the corrector performs one step per predictor iteration, and a total of 30 predictor steps per diffusion block.
1
Input : FOA ${\boldsymbol{A}}^{\tau}_{1}=[{\boldsymbol{a}}_{1}(1,\ldots,{\tau_{\max}})]\in\mathbb{C}^{4\times{\tau_{\max}}}$
Score models ${\boldsymbol{s}}_{1}(\;\cdot\;;{\boldsymbol{\theta}}_{1})$, ${\boldsymbol{s}}_{2}(\;\cdot\;;{\boldsymbol{\theta}}_{2})$
2 for $N=1$ to 22 do
3  Set ${\boldsymbol{A}}^{f}_{N}\leftarrow{\rm STFT}({\boldsymbol{A}}^{\tau}_{N})$ ;
4   Transform amplitude via ${\boldsymbol{y}}\leftarrow\mathcal{R}\{\mathcal{H}({\boldsymbol{A}}^{f}_{N})\}$;
5   Generate $2(2N+3)$ channels ${\boldsymbol{x}}_{T}\sim\mathcal{N}({\boldsymbol{0}},\sigma_{\text{max}}{\boldsymbol{I}})$;
6   Sample $2(2N+3)$ conditioned channels as ${\boldsymbol{x}}_{0}={\rm PC\_Sample}({\boldsymbol{x}}_{T},{\boldsymbol{s}}_{N}(\;\cdot\;,{\boldsymbol{y}};\theta_{N}))$;
7   Set AU ${\boldsymbol{A}}^{f}_{N+1}\leftarrow[{\boldsymbol{A}}^{f}_{N},\mathcal{H}^{-1}(\mathcal{C}({\boldsymbol{x}}_{0}))]$;
8   Compute ${\boldsymbol{A}}^{\tau}_{N+1}\leftarrow{\rm ISTFT}({\boldsymbol{A}}^{f}_{N+1})$
return HOA ${\boldsymbol{A}}^{\tau}_{3}\in\mathbb{C}^{16\times{\tau_{\max}}}$
Algorithm 1 DiffAU
3.2 Training
Algorithm 1 requires trained SGMs $\{{\boldsymbol{\theta}}_{N}\}_{N=1}^{2}$. In DiffAU, each diffusion block is trained independently, and the blocks are combined afterwards. The training follows the denoising score matching strategy [22], where the model is optimized to estimate the gradient of the log probability of noisy data, i.e., the log of $p_{t}$, which represents the distribution of ${\boldsymbol{x}}_{0}+\sigma_{t}\cdot{\boldsymbol{z}}$ with ${\boldsymbol{x}}_{0}\sim p_{\rm data}$ and ${\boldsymbol{z}}\sim\mathcal{N}(0,{\boldsymbol{I}})$ [23]. This approach leverages the Markovian structure of the forward diffusion process (3) and the Gaussian noise, allowing ${\boldsymbol{x}}_{t}$ to be sampled in a single step..
Under this framework, the training method effectively learns a separate denoiser for each timestep $t$ by predicting the noise. To construct a dataset for upscaling to order-$N+1$, each data pair consists of an order-$N$ Ambisonics signal $\boldsymbol{y}$}, and the corresponding ${\boldsymbol{x}}_{0}\in\mathbb{R}^{2N+3}$, which contains the additional $2N+3$ channels required to reach order-$N+1$. The resulting dataset is $\mathcal{D}=\{({\boldsymbol{x}}_{0}^{(i)},{\boldsymbol{y}}^{(i)})\}_{i=0}^{D}$, and the empirical risk based on $\mathcal{D}$ is given by
$\hskip-5.69046pt\mathcal{L}_{\mathcal{D}}({\boldsymbol{\theta}}_{N})\!=\!\frac{1}{D}\sum_{i=1}^{D}\left\|s_{N}\big({\boldsymbol{x}}_{t^{(i)}}^{(i)},{\boldsymbol{y}}^{(i)};{\boldsymbol{\theta}}_{N}\theta)\cdot\sigma_{t^{{}^{(i)}}}\!+\!{\boldsymbol{z}}^{(i)}\right\|^{2},$ (7)
where $t^{(i)}\sim\mathcal{U}[1,T]$ and ${\boldsymbol{z}}^{(i)}\sim\mathcal{N}(0,{\boldsymbol{I}})$.
3.3 Discussion
DiffAU leverages the generative capabilities of diffusion models to sample from the posterior distribution, enabling physically plausible solutions to the ill-posed AU. It adapts diffusion models to match the Ambisonics signal format and employs a cascaded structure to estimated the HOA channels. The cascaded architecture introduces modularity, allowing upscaling from any desired order. While we consider noiseless, multiple-speaker scenarios in free-field conditions, we expect its design to yield accurate AU also under noise and reverberation, while leaving this gextension for future work.
4 Numerical Study
We evaluate the proposed DiffAU in a numerical study [^1] The source code and the complete set of hyperparameters used in our study is available at https://github.com/Amitmils/DiffAU. detailed next, and in a listening experiment detailed in Section 5.
Data: We constructed a dataset based on the WSJ0 corpus [24]. The dataset is split into training, validation, and test sets, with each speaker appearing in only one set and contributes multiple utterances to it. To generate the Ambisonic signal sets, we randomly selected 141-4 speakers for each signal. For each selected speaker, a random DOA was assigned, and the corresponding 3rd order Ambisonics was constructed via (1). This study focuses exclusively on free-field scenarios. All audio signals are 2.048 seconds long and sampled at 16 kHz.
Evaluation: Our evaluation uses the FOA (first four channels) as input. The model then estimates the remaining twelve channels to reconstruct the HOA. These predicted channels are compared to the corresponding ground truth using the STFT signal-to-distortion ratio (STFT-SDR) metric, computed over all higher-order channels (i.e., channels 5–16) for each sample:
$\text{STFT-SDR}(\hat{{\boldsymbol{A}}}^{\prime}_{3})\!=\!10\cdot\log_{10}\left(\frac{\|{\boldsymbol{A}}^{f}_{3_{(5:16)}}\|_{F}^{2}}{\|{\boldsymbol{A}}^{f}_{3_{(5:16)}}\!-\!\hat{{\boldsymbol{A}}}^{f}_{3_{(5:16)}}\|_{F}^{2}}\right)$ (8)
where ${\boldsymbol{A}}^{f}_{3}$ and $\hat{{\boldsymbol{A}}}^{f}_{3_{(5:16)}}$ are the true and estimated Ambisonic signals, respectively, in the TF domain. The subscript (5:16){(5:16)} indicates the channels used, and $\|\cdot\|_{F}$ is the Frobenius norm.
Results: The STFT-SDR results in Table 1 are based on 500 audio samples, corresponding to 0.25 hours of test data. Evaluation focuses on the HOA channels (channels 5–16). DiffAU was trained with 10 hours of data per diffusion block. We compare its performance to the PWD method using CS in the frequency domain [5], which is an iterative method that addresses the underdetermined nature of AU by imposing sparsity on the sound field. Table 1 clearly shows that DiffAU outperforms the baseline for all cases. Although we consider a free-field setting, where the sparsity assumption holds, DiffAU outperforms PWD CS by directly learning the posterior.
In Fig. 2 compares directional energy plots of 2nd- and 3rd-order Ambisonics signals produced by DiffAU with the ground truth for orders 1–3, across one to four active speakers. The results show a strong resemblance between the recovered HOA and the reference energy patterns.
Table 1: STFT-SDR results in dB on the HOA channels
# Speakers DiffAU PWD CS # Audios
1 $29.5\pm 6.7$ $2.9\pm 7.9 11$5
2 $27.3\pm 3.8$ $14.3\pm 2.2$ 127
3 $23.1\pm 4.0$ $12.3\pm 2.6$ 131
4 $19.6\pm 4.5$ $10.9\pm 2.6$ 127
Overall $\mathbf{24.7}\pm\mathbf{6.2}$ $12.6\pm 4.5$ 500
5 Listening Test
Informal listening in the previous section suggested that signals estimated by both the AU and PWD CS methods were largely indistinguishable from the reference, consistent with their low reconstruction errors. The formal listening test aimed to assess whether DiffAU introduced subtle audible artifacts potentially undetectable by error metrics.
Fig. 2: Directional energy plots (azimuth-elevation). Columns: FOA, 2nd- and 3rd-order Ambisonics ground truth, 2nd- and 3rd-order DiffAU outputs. Rows correspond to the number of active sources.
Following the MUltiple Stimuli with Hidden Reference and Anchor (MUSHRA) protocol [25], participants evaluated overall signal quality relative to a reference across three MUSHRA screens. Each screen presented three signals: the 3rd-order Ambisonics reference, a 1st-order Ambisonics anchor, and the DiffAU-upscaled 3rd-order signal. All signals were rendered binaurally using the least-squares method [26] and loudness-equalized to –6 LUFS; no head-tracking was applied. Each screen featured a single speaker from the test set at a colatitude of $90^{\circ}$ and azimuths of $15^{\circ}$, $30^{\circ}$, and $60^{\circ}$. Nine participants (6 male, 3 female) with prior spatial listening experience and no known hearing impairments took part. The test was conducted in a quiet environment and included training and familiarization phases. During training, participants were introduced to the equipment and MUSHRA scale, while familiarization allowed free listening to all stimuli. Participants then rated the perceptual quality of each signal on a 0–100 scale, where 100 indicates no audible difference from the reference.
Fig. 3: Listening test results
As shown in Fig.3, our method was perceptually indistinguishable from the 3rd-order Ambisonics reference, which received a perfect score of 100 from all participants. Our method scored 100 on all screens except for one participant, who rated two screens 94 and 95. The 1st-order signal averaged 28.7. These results align with the high STFT-SDR reported in Section4, indicating perceptual equivalence with no transient artifacts. We note, however, that this outcome may reflect the free-field evaluation, and performance in more realistic environments remains to be tested.
6 Conclusion
We proposed a novel AU method termed DiffAU. By leveraging diffusion models to sample from the posterior distribution, DiffAU addresses the inherent underdetermined nature of the AU problem. For multi-speaker scenarios in free-field conditions, DiffAU outperforms the baseline, and was indistinguishable from true HOA from a perceptual standpoint.